OpenStudy (anonymous):

Solve using either substitution or elimination process for the nonlinear equations: x^2+y^2=36 and x+y=6

OpenStudy (anonymous):

do i use substitution or elimination?

OpenStudy (anonymous):

\[x^2+y^2=36\] \[x+y=6\] \[y=6-x\] \[x^2+(6-x)^2=36\]

OpenStudy (anonymous):

that would be "substitution"

OpenStudy (anonymous):

If one of the equations given to you is very simple to put in a (for example) "x=" or "y=" form, use substitution.

OpenStudy (anonymous):

so how do i take this info and find the solutions

OpenStudy (anonymous):

by solving the quadratic equation \[x^2+(6-x)^2=36\]

OpenStudy (anonymous):

\[x ^{2}+(6-x)^{2}=36 \] meaning i would subtract 36?

OpenStudy (anonymous):

\[x^2+(6-x)^2=36\] \[x^2+(6-x)(6-x)=36\] \[x^2+36-12x+x^2=36\] \[2x^2-12x=0\] \[2x(x-6)=0\] \[x=0,x=6\]

OpenStudy (anonymous):

omg thank you :)

OpenStudy (anonymous):

so i would have 3 real solutions for this equation?

OpenStudy (anonymous):

no just 2

OpenStudy (anonymous):

0 and 6

OpenStudy (anonymous):

okay because y='s that equation which isnt an answer

OpenStudy (anonymous):

if x = 0, since \[x+y=6\] that means \[0+y=6\implies y =6\]

OpenStudy (anonymous):

and if x = 6 we get \[6+y=6\implies y=0\]

OpenStudy (anonymous):

right the answers are written: Two real solutions \[(\pm6, 0) \And (0,\pm6) \]

OpenStudy (anonymous):

remove the plus and minus sign.